Logiweb bibliography of check

Up Help

[0] check

[1] base

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-16.UTC:14:58:03.474266 = MJD-53902.TAI:14:58:36.474266 = LGT-4657186716474266e-6